Ninety Passengers rode in a train from City A to City B. Tickets for regular coach seats cost $121. Tickets for sleeper seats cost $281. The receipts for the trip totaled $18,090. How many passengers purchased each type of ticket?
Answer by Boreal(15235)

C+S=90
121C+281S=18,090
-121C-121S=-10,890
160S=7200
S=45@281=$12,645
C=45@121=$5445
They add to $18,090.
45 passengers for each.
(I'm a volunteer tutor).